OpenStudy (anonymous):

x^2+10x+16 x+8 ______________ / ________ x^2-6x-16 x^2-64

OpenStudy (anonymous):

so i need help factoring this equation

OpenStudy (bibby):

first things first rewrite it so that it's multiplied division = multiplication by the reciprocal

OpenStudy (anonymous):

so do the keep change flip thing?

OpenStudy (bibby):

for example dividing by 3/2 is the same t hing as multiplying by 2/3

OpenStudy (bibby):

I don't remember the keep change thing

OpenStudy (anonymous):

wait were u talking about changing it to multiplication?

OpenStudy (bibby):

yeah. \(\large \frac{x^2+10x+16 }{x^2-6x-16 }\div \frac{x+8}{x^2-64}=\frac{x^2+10x+16 }{x^2-6x-16 }*\frac{x^2-64}{x+8}\)

OpenStudy (anonymous):

yeah thats what i was talking about

OpenStudy (bibby):

nowe we start factoring and canceling

OpenStudy (anonymous):

yeah thats what i have a hard time with the factoring

OpenStudy (bibby):

same

OpenStudy (bibby):

factor \((x^2-64)\)

OpenStudy (bibby):

this might help you out \((a^2-b^2)=(a+b)(a-b)\)

OpenStudy (anonymous):

that be like (x+8)(x-8) right?

OpenStudy (bibby):

yas

OpenStudy (bibby):

what about \(x^2+10x+16\)

OpenStudy (anonymous):

(x+8)(x+2) ?

OpenStudy (bibby):

correct. so let's update our fraction \(\huge\frac{(x+8)(x+2)}{x^2-6x-16 }*\frac{(x+8)(x-8)}{x+8}\)

OpenStudy (anonymous):

ok so now we need x^2-6x-16 but im having trouble with that one

OpenStudy (bibby):

list out the factors of 16

OpenStudy (bibby):

1,2,4,8,16 what combination of these can we use to get -6?

OpenStudy (anonymous):

2 and 4? but that doesn't multiply into 16

OpenStudy (bibby):

8 and 2 the way it's written the pairs match outwards 1-16 2-8 4-4

OpenStudy (anonymous):

so it would be (x+8)(X-2)?

OpenStudy (bibby):

that would be positive 6, we want -6 the way I think about it is 2-8=-6 or -8+2=-6

OpenStudy (anonymous):

ohh oops ok so (x-8)(x+2)

OpenStudy (bibby):

looks gewd

OpenStudy (bibby):

time to reupdate the fraction

OpenStudy (bibby):

\[\huge\frac{(x+8)(x+2)}{(x-8)(x+2) }*\frac{(x+8)(x-8)}{x+8}\]

OpenStudy (anonymous):

so i can cross out (x+2) right?

OpenStudy (bibby):

yeah

OpenStudy (anonymous):

which (x+8 do i keep?

OpenStudy (bibby):

whichever you want. conceptually, it's one long fraction

OpenStudy (anonymous):

does it matter actually? that should be my answer right?

OpenStudy (bibby):

\(\huge\frac{(x+8)\cancel{(x+2)}}{(x-8)\cancel{(x+2)} }*\frac{(x+8)(x-8)}{x+8}\) \(\huge\frac{(x+8)(x+8)(x-8)}{(x-8) (x+8)}{}\)

OpenStudy (bibby):

then we can cancel an x+8 and x-8

OpenStudy (anonymous):

so yeah the answer would just be x+8

OpenStudy (bibby):

yep

OpenStudy (anonymous):

can u help me with another one?

OpenStudy (bibby):

sure

OpenStudy (anonymous):

\[\frac{ 6x^2-32x+10 }{ 3x^2-15x}\div \frac{ 3x^2+11x-4 }{ 2x^2-32 }\]

OpenStudy (bibby):

ew I never learned how to do the ones with a>1

OpenStudy (bibby):

let me google it

OpenStudy (bibby):

for starters we can factor \(2x^2-32\)

OpenStudy (bibby):

factor 2 out. \(2(x^2-16)\)

OpenStudy (anonymous):

2(x+8)(x-8)?

OpenStudy (bibby):

4*

OpenStudy (bibby):

16=4^2

OpenStudy (anonymous):

+4 or -4?

OpenStudy (bibby):

both

OpenStudy (bibby):

a^2-b^2=(a+b)(a-b) x^2-16=x^2-4^2

OpenStudy (anonymous):

what? lol how would u write it?

OpenStudy (bibby):

\(a^2-b^2=(a+b)(a-b)\) \(x^2-4^2=x+4(x-4)\)

OpenStudy (anonymous):

wait but then where would the 2 go?

OpenStudy (bibby):

what we have is 2(x+4)(x-4) we can either leave it like that or write it as \(((2*x)+(2*4))(x-4)=(2x+8)(x-4)\)

OpenStudy (anonymous):

OHHHHHHHHHHH!!!!!!

OpenStudy (bibby):

woah, forgot to google the other thing, give me 2 minutes

OpenStudy (anonymous):

okieieee

OpenStudy (bibby):

\(ax^2+bx+c=0\) Step 1: Find two numbers that multiply to give ac (in other words a times c), and add to give b. step 2: rewrite the middle

OpenStudy (bibby):

\(6x^2-32x+10\) ac=6*10=60 b=-32 not sure how that works out here let's do the thing we did before with \(3x^2-15x\)

OpenStudy (anonymous):

ok that be 3(x-5) right?

OpenStudy (bibby):

I'd factor out 3x

OpenStudy (anonymous):

how?

OpenStudy (bibby):

\(3x(\frac{3x^2}{3x}-\frac{15x}{3x})\)

OpenStudy (bibby):

so that comes out to 3x(x-5)

OpenStudy (anonymous):

ohhhhh i for got the x

OpenStudy (bibby):

yus

OpenStudy (bibby):

uhh let's see where we're at

OpenStudy (bibby):

\(\huge \frac{ 6x^2-32x+10 }{ (3x)(x-5)}\times\frac{2(x+4)(x-4) }{ 3x^2+11x-4 }\) I'm going to cheat for the other 2

OpenStudy (anonymous):

okayy

OpenStudy (bibby):

\(6x^2-32x+10=2 (x-5) (3 x-1)\)

OpenStudy (bibby):

\( 3x^2+11x-4=(3 x-1) (x+4)\)

OpenStudy (anonymous):

so x-5 crosses out and so does x+4

OpenStudy (bibby):

btw I'm using this http://www.wolframalpha.com/input/?i=factor+3x%5E2%2B11x-4 \(\huge \frac{ (x-5)(3x-1) }{ (3x)(x-5)}\times\frac{2(x+4)(x-4) }{ (3x-1)(x+4) }\)

OpenStudy (anonymous):

oh an so does 3x-1

OpenStudy (bibby):

\(\huge \frac{ \cancel{(x-5)}\cancel{(3x-1)} }{ (3x)\cancel{(x-5)}}\times\frac{2\cancel{(x+4)}(x-4) }{ \cancel{(3x-1)}\cancel{(x+4)} }\) looks like the canceling leaves us with \(\huge \frac{2(x-4)}{3x}\)

OpenStudy (anonymous):

that website helps factor?

OpenStudy (bibby):

it does everything. calculus, probability whatever you say "factor XXXXXXXXXXXXXX" it factors you say "derive YYYYYYYYYYYY" it derives

OpenStudy (anonymous):

:o this is amazing

OpenStudy (anonymous):

do u have to factor with subtraction and addition?

OpenStudy (bibby):

I know! I'm sure there are other calculators online that do this too but if you want my honest opinion, you should check out a factoring guide on youtube what do you m ean

OpenStudy (anonymous):

like for this... \[\frac{ 6y-4 }{ y^2-5}+\frac{ 3y+1 }{ y^2-5 }\]

OpenStudy (bibby):

it should work you just have to learn how to use parentheses. watch ``` simplify ((6y-4)/(y^2-5))+((3y+1)/(y^2-5)) ``` it's messy http://www.wolframalpha.com/input/?i=simplify+%28%286y-4%29%2F%28y%5E2-5%29%29%2B%28%283y%2B1%29%2F%28y%5E2-5%29%29 btw that's in common base so you can just rewrite it as \(\large \frac{6y-4+(3y+1)}{y^2-5}\)

OpenStudy (anonymous):

is that the answer? do u combine like terms?

OpenStudy (bibby):

yeah, always

OpenStudy (bibby):

9y-3

OpenStudy (anonymous):

i see wb this one?

OpenStudy (bibby):

dammm bby getting a bit out of hand

OpenStudy (bibby):

jk not doing anything of value right now anyhow

OpenStudy (anonymous):

\[\frac{ 3 }{ 8x^3y^3 }-\frac{ 1 }{ 4xy }\]

OpenStudy (anonymous):

lol sorry after this one i have like 2 more questions so im like merp wanna finish

OpenStudy (bibby):

factor the bottom left denominator

OpenStudy (bibby):

yea, we'll finish

OpenStudy (anonymous):

how do i do that

OpenStudy (bibby):

there are 2 ways of thinking about it. we want the least common multiply of 4xy and 8x^3y^3 what my brain instinctively says is factor 4 x and y \(8x^3y^3=4xy(2x^2y^2)\)

OpenStudy (bibby):

then we know that if we multiply the right hand side by \(\large \frac{2x^2y^2}{2x^2y^2}\) (which is effectively 1, we'll have a common denominator.

OpenStudy (anonymous):

so the bottom right side would turn into 8x^3y^3 and the top would be 2X^2y^2?

OpenStudy (bibby):

yeah